OpenStudy (anonymous):

please help me!! tell weather the sequence is arithmetic. if it is what is the common difference.2,7,13,20 A)yes 5 B) yes 6 C) yes 2 D)no

OpenStudy (shinalcantara):

try to examine your given data is 20-13 equal to 13-7 & 7-2?

OpenStudy (anonymous):

7-2= 5 13-7=6 20-13= 7 Since the difference between any two consecutive numbers is not the same hence the given sequence is not an arithmetic sequence. D)no

OpenStudy (ilovecake):

@dpasingh is correct

OpenStudy (anonymous):

Thank you so much:)